OpenStudy (anonymous):

Calculate the derivatives of the following real functions by using the chain rule; maybe several times. In each case, determine the inner and outer function to be used, together with their respective areas of definition. b) \(h(x) = exp(\frac{1}{sin x})\)

OpenStudy (kinggeorge):

First, I would write it as \[\large h(x)=e^{\frac{1}{\sin(x)}}\]Now, can you tell me what the outer function is?

OpenStudy (anonymous):

i have no idea to be honest, its also important for mx exam, which part of math should i learn, can you send me maybe a link George?

OpenStudy (anonymous):

i know i must memorize some stuff for derivatives..

OpenStudy (kinggeorge):

Looking at this, you can tell that the outer function is \(e^u\) if you let \(u=\frac{1}{\sin(x)}\). Recall that the derivative of \(e^u\) is \(e^u\) itself. That means you're looking for \[\large \frac{d}{dx}\;e^{\frac{1}{\sin(x)}}=e^{\frac{1}{\sin(x)}}\cdot \frac{d}{dx}\;\frac{1}{\sin(x)}\]

OpenStudy (anonymous):

ok..

OpenStudy (kinggeorge):

Now rewrite \[\frac{1}{\sin(x)}=\sin^{-1}(x).\]We now apply the chain rule again, this time using the power rule to differentiate. We get \[\large e^{\frac{1}{\sin(x)}}\cdot \frac{d}{dx}\;\sin^{-1}(x)=e^{\frac{1}{\sin(x)}}\cdot -\sin^{-2}(x)\cdot \frac{d}{dx}\;\sin(x)\]

OpenStudy (kinggeorge):

We have one more step to go. We just need to take the derivative of \(\sin(x)\). That means the final derivative should be \[\large e^{\sin^{-1}(x)}\cdot -\sin^{-2}(x)\cdot \frac{d}{dx}\;\sin(x)= e^{\sin^{-1}(x)}\cdot -\sin^{-2}(x)\cdot\cos(x)\]We can rewrite this in fraction form to be \[\large e^{\frac{1}{\sin(x)}}\cdot -\frac{\cos(x)}{\sin^{2}(x)}\]

OpenStudy (anonymous):

ok there is a lot ofs thing to memorize learn for me i think, how can i search for this thema what is the name of this topic ?

OpenStudy (kinggeorge):

In general, the best way to memorize these is to do them. However, there is some useful information here http://en.wikipedia.org/wiki/Table_of_derivatives and here http://en.wikipedia.org/wiki/Derivative#Computing_the_derivative

OpenStudy (anonymous):

ok this is exact what i want, i need to make me warm slowly for exam..

OpenStudy (kinggeorge):

We're not done with the problem yet however. We still need to list the functions we looked at, and they're respective domains. The first function (the outer function) is just \(e^x\). This has a domain of all reals. The next is the function \(\frac{1}{x}\). This has a domain of all reals except 0. The final function was \(\sin(x)\) which has a domain of all reals. Only the first function is an outer function. The rest are inner functions.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

pls feel free to go on

OpenStudy (kinggeorge):

That was the rest of it right there.

OpenStudy (anonymous):

aah ok thank you George :) sorry

OpenStudy (kinggeorge):

You're welcome.